x dy/dx = y + x


1
Expert's answer
2021-02-24T05:56:29-0500

"xy'=y+x"

"y'=\frac{y}{x}+1"

Let "y=zx", then "\frac{y}{x}=z" and "y'=z+z'x". Then we have

"z+z'x=z+1", so "z'=\frac{1}{x}"

"z=\ln|x|+C", so we have "y=zx=x\ln|x|+Cx"

Answer: "y=x\ln|x|+Cx"